# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 2. Notations and statement of the result
where $\Delta(e^{2i\pi\theta})=2^{n(n-1)/2}\prod_{i<j}\sin\left(\pi(\theta_{i}-\theta_{j})\right)$ for $\theta\in\mathcal{H}$, $Vol_{g}$ denotes the volume with respect to the Lebesgue measure on $\mathbb{R}^{Supp(g)\setminus\partial R_{d,n}}$.
Note that the case $n=2$ admits explicit formulas which do not need such a machinery. Some polytopes $P^{g}_{\alpha,\beta,\gamma}$ may be empty and thus may not contribute to $d\mathbb{P}[\gamma|\alpha,\beta]$. Numerical experiments for $n=3$ suggest that there are $\alpha,\beta\in\mathcal{H}_{reg}$ for which all regular labeling $g$ yields a non-empty polytope $P^{g}_{\alpha,\beta,\gamma}$ for some $\gamma\in\mathcal{H}_{reg}$. However, for a fixed triple $(\alpha,\beta,\gamma)\in\mathcal{H}_{reg}$ there seems to be generically only a strict subset of $\{P^{g}_{\alpha,\beta,\gamma}\}_{g\text{ regular}}$ which are not empty and contribute. Finally, remark that we only considered the case of regular conjugacy classes to ensure the existence of a density for the convolution product. Such a hypothesis is regularly assumed (see for example [31, 23]). However, we expect similar results to hold in cases where less than $n/2$ coordinates of $\mathcal{H}$ are equal, in which case the convolution product is still expected to have a density.
As a direct application of this theorem, we deduce a positive formula for the volume of flat connections on the punctured two-dimensional sphere with prescribed holonomies around the holes. Denote by $\Sigma_{0}^{3}$ the sphere with three removed marked points $a,b,c$. We denote by $\mathcal{M}(\Sigma_{0}^{3},\alpha,\beta,\gamma)$ the moduli space of flat $SU_{n}$-valued connections on $\Sigma_{0}^{3}$ for which the holonomies around $a,b,c$ respectively belongs to $\mathcal{O}_{\alpha},\mathcal{O}_{\beta}$ and $\mathcal{O}_{\gamma}$. In the specific case of the punctured sphere, this moduli space can be alternatively described as
@@
where $SU_{n}$ acts diagonally by conjugation, see [1], see also [23, Section 3] for an introduction to $SU_{n}$-valued flat connections and a proof of the latter equality. As a corollary of Theorem 2.4, we thus get an expression of the volume of $\mathcal{M}(\Sigma_{0}^{3},\alpha,\beta,\gamma)$ as a sum of volumes of explicit polytopes.
Corollary 2.5.
Let $n\geq 3$ and consider the usual volume on $SU_{n}$. For $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ such that $|\alpha|_{1},\,|\beta|_{1},\,|\gamma|_{1}\in\mathbb{N}$, then $Vol\left[(\mathcal{M}(\Sigma_{0}^{3},\alpha,\beta,\gamma)\right]\not=0$ only if $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}+\sum_{i=1}^{d}\gamma_{i}=n+d$ for some $d\in\mathbb{N}$, in which case
@@
where $\widetilde{\gamma}=(1-\gamma_{n},\ldots,1-\gamma_{1})$ and the polytope $P^{g}_{\alpha,\beta,\widetilde{\gamma}}$ are defined in Theorem 2.4.
Note that the choice of normalization for the volume of $SU_{n}$ slightly differs from the one used in [31] for numerical applications. A consequence of this corollary is that the volume is a piecewise polynomial in $\alpha,\beta,\gamma$, up to the normalization factor coming from the volume of the conjugacy classes. Such a phenomenon is a reflect of the underlying symplectic structure which had already been observed in [23]. A same phenomenon occurs in the co-adjoint case, see [7, 10]. Remark that the results of Corollary 2.5 only address the genus zero case when the Lie group is $SU_{n}$. It would be very interesting to have similar expressions in higher genus cases or for different compact Lie groups. An equally interesting problem would be to get such explicit expressions already at the level of the fusion coefficients.
@@
Let us give the structure of the paper while sketching the proof of our main statement. The first step of the prood is the semi-classical approximation of the density of the convolution product by a limit of quantum Littlewood-Richardson coefficients. Such approximation scheme is done in Section 3. All the work of the remaining part of the manuscript consists in turning known expressions for the quantum Littlewood-Richardson coefficients into integers points counting in convex bodies, for which the convergence towards volumes of polytope is straightforward, see [16]. Section 4 introduces the puzzle expression for those coefficients obtained in [5] and gives a first simplification of the puzzle formulation by only keeping the position of certain pieces of the puzzles. It is then deduced in Section 5 an expression of the coefficients as the counting of integers points in a family of convex polytopes indexed by certain two-colored tilings which are reminiscent of Figure 4. Up to this point, those polytopes are degenerated and non-rational polytopes in a higher dimension space, which prevents any proper asymptotic counting in the semi-classical limit. By a combinatorial work on the underlying two-colored tiling, we give in Section 6 a parametrization of the integer points of those polytopes in terms of integer points of genuine convex bodies. Remark that the results of Section 5 hold more generally for any coefficient of the two-step flag variety, a fact which is not true anymore from Section 6. The asymptotic counting of integers points in convex bodies is then much more tractable, and the conclusion of the proof of Theorem 2.4 and Corollary 2.5 is done in Section 7.
3. Density formula via the quantum cohomology of the Grassmannian
Let $n\geq 1$ and consider $\alpha,\beta\in\mathcal{H}^{2}$ : $\alpha=(\alpha_{1},\dots,\alpha_{n})$, $\beta=(\beta_{1},\dots,\beta_{n})$ where
@@ 3. Density formula via the quantum cohomology of the Grassmannian
Suppose furthermore that
@@
for some $k,k^{\prime}\in\mathbb{Z}$. Let $A=U\mathrm{e}^{2i\pi\alpha}U^{*},B=V\mathrm{e}^{2i\pi\beta}V^{*}$, where $U,V$ are independent Haar distributed matrices on $U(n)$ so that $A$ and $B$ are respectively uniformly distributed on the conjugacy classes $\mathcal{O}(\alpha)$ and $\mathcal{O}(\beta)$. The goal of this section is to prove Theorem 3.8 stated below which gives a density formula (3.28) for the probability $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ that $AB\in\mathcal{O}(\gamma)$ for $\gamma\in\mathcal{H}$ in terms of structure constants defined in Section 3.2. Such a semi-classical convergence had been already suggested and proven several times in different forms (see [31] for a similar approach with fusion coefficient and [8] for a convergence in distribution). In Section 3.1, we recall in Proposition 3.1 a classical expression of the density in terms of characters of irreducible representations of $SU(n)$. In Section 3.2, we link the density of Proposition 3.1 to the structure constants and derive Theorem 3.8.
@@ 3.1. A first density formula
This part aims at recalling a proof of the formula (3.3) which gives the value of $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ as an infinite sum of characters. A similar treatment of the convolution of orbit measures in the general context of Lie algebras can be found in [8, Sec. 7].
@@
the push-forward of $\mathrm{d}g$ by $\varphi_{\theta}$. The measure $m_{\theta}$ is a measure on $\mathcal{O}(\theta)$ called the orbital measure. For any function $f:\mathcal{O}(\theta)\rightarrow\mathbb{R}$,
@@
Recall that the irreducible representations of the compact group $SU(n)$ are parameterized by $\lambda\in\mathbb{Z}_{\geq 0}^{n}$ and we denote by $(\rho_{\lambda},V_{\lambda})$ the corresponding representation where $\rho_{\lambda}:SU(n)\rightarrow V_{\lambda}$ and $\chi_{\lambda}:End_{V_{\lambda}}\rightarrow\mathbb{C},x\mapsto Tr[x]$ is the associated character.
@@
Let $(\alpha,\beta)\in\mathcal{H}^{2}$ and let $A,B\in\mathcal{O}(\alpha)\times\mathcal{O}(\beta)$ be two independent random variables sampled from $m_{\alpha}$ and $m_{\beta}$ respectively. Let $C=AB\in\mathcal{O}(\gamma)$ for some random $\gamma\in[0,1[^{n}/S_{n}$. The density of $\gamma=\gamma_{1}>\dots>\gamma_{n}\geq 0$ is given by the absolute convergent series
(3.3) $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]=\frac{|\Delta(\mathrm{e}^{2i\pi\gamma})|^{2}}{(2\pi)^{n-1}n!}\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\frac{1}{\dim V_{\lambda}}\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\chi_{\lambda}(\mathrm{e}^{-2i\pi\gamma}).$
Another expression of the density (3.3) is given in (3.26).
@@
(3.4) $\widehat{m}:\mathbb{Z}_{\geq 0}^{n}$ $\rightarrow\mathrm{End}_{V_{\lambda}}$
(3.5) $\lambda$ $\mapsto\int_{SU(n)}\rho_{\lambda}(g)\mathrm{d}m(g).$
From this one can deduce the Fourier transform of $m_{\theta}$.
@@
For any $\lambda\in\mathbb{Z}_{\geq 0}^{n}$, the function $g\mapsto\rho_{\lambda}(g)$ is invariant by conjugation, so that
@@
satisfies $\widehat{m_{\theta}}(\lambda)=c\cdot\mathrm{id}_{V_{\lambda}}$. One computes the value of $c$ by taking the trace which gives
@@
Let $m,m^{\prime}$ be two measures on $SU(n)$. Let $m\boxtimes m^{\prime}$ be the product measure on $SU(n)\times SU(n)$. Define $mult:SU(n)\times SU(n)\rightarrow SU(n)$ to be the multiplication on $SU(n)$ : $mult(g_{1},g_{2})=g_{1}g_{2}$. The convolution of $m$ and $m^{\prime}$ denote by $m*m^{\prime}$ is defined as
(3.9) $m*m^{\prime}=mult\#(m\boxtimes m^{\prime}).$
@@
By Definition 3.4, the measure $m_{\alpha,\beta}$ is the law on $SU(n)$ of $C=A\cdot B$ where $A$ and $B$ are sampled from measures $\mu_{\alpha}$ and $\mu_{\beta}$ on $\mathcal{O}(\alpha)$ and $\mathcal{O}(\beta)$ respectively. Recall that for two measures $m,m^{\prime}$ on $SU(n)$, then,
@@
Let $f:\mathbb{Z}_{\geq 0}^{n}\rightarrow End_{V_{\lambda}}$ be a function such that
(3.12) $\left\lVert f\right\rVert^{2}=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\dim V_{\lambda}\cdot Tr[f(\lambda)f^{*}(\lambda)]<\infty.$
@@
(3.14) $g$ $\mapsto\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\dim V_{\lambda}\cdot Tr[\rho_{\lambda}(g^{-1})f(\lambda)].$
@@
For $(\alpha,\beta)\in\mathcal{H}^{2}$,
(3.15) $\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\dim(V_{\lambda})Tr[\widehat{m}_{\alpha,\beta}(\lambda)\widehat{m}_{\alpha,\beta}(\lambda)^{*}]<\infty.$
@@
(3.16) $\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\dim(V_{\lambda})Tr[\widehat{m}_{\alpha,\beta}(\lambda)\widehat{m}_{\alpha,\beta}(\lambda)^{*}]=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\frac{1}{\dim(V_{\lambda})^{2}}\left|\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\right|^{2}\left|\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\right|^{2}.$
Using Weyl’s character formula [7, eq. (21)],
@@
where $\lambda^{\prime}=\lambda+\rho$ with with $\rho=(n-1,\ldots,0)$ and where $\Delta(x_{1},\dots,x_{n})=\prod_{1\leq i<j\leq n}(x_{i}-x_{j})$ is the Vandermonde determinant. Recall that by assumption, $\alpha_{i}\neq\alpha_{j}$ for $i\neq j$ and the same holds for $\beta$, so that the expressions $\Delta(\mathrm{e}^{2i\pi\alpha})$ and $\Delta(\mathrm{e}^{2i\pi\beta})$ are well-defined. The previous sum becomes
@@
which can be found in [7, eq. (19)] and where $sf(n)=\prod_{1\leq j\leq n}j!$, it suffices to show that
(3.21) $V_{n}=\sum_{\lambda_{1}>\dots>\lambda_{n}>0}\frac{1}{\Delta(\lambda)^{2}}$
converges. Let us write
$\sum_{\lambda_{1}>\dots>\lambda_{n}>0}\frac{1}{\Delta(\lambda)^{2}}$ $=\sum_{\lambda_{1}>\dots>\lambda_{n}>0}\prod_{1\leq i<j\leq n}(\lambda_{i}-\lambda_{j})^{-2}$
$=\sum_{\lambda_{2}>\dots>\lambda_{n}>0}\left[\sum_{\lambda_{1}>\lambda_{2}}\prod_{2\leq j\leq n}(\lambda_{1}-\lambda_{j})^{-2}\right]\prod_{2\leq i<j\leq n}(\lambda_{i}-\lambda_{j})^{-2}$
$\leq\sum_{\lambda_{2}>\dots>\lambda_{n}>0}\left[\sum_{\lambda_{1}>\lambda_{2}}(\lambda_{1}-\lambda_{2})^{-2(n-1)}\right]\prod_{2\leq i<j\leq n}(\lambda_{i}-\lambda_{j})^{-2}$
the innermost sum is bounded by $\sum_{k\geq 1}k^{-2(n-1)}\leq\sum_{k\geq 1}k^{-2}=c$ for $n\geq 2$. Thus,
$V_{n}\leq cV_{n-1}$
so that $V_{n}\leq c^{n-1}$ which proves the convergence.
@@
Let $(\alpha,\beta)\in\mathcal{H}^{2}$ and $g\in SU(n)$. Then,
(3.22) $(\widehat{m}_{\alpha,\beta})^{\lor}(g)=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\frac{1}{\dim V_{\lambda}}\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\chi_{\lambda}(g^{-1})$
@@
(3.23) $(\widehat{m}_{\alpha,\beta})^{\lor}(g)$ $=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\dim V_{\lambda}\cdot Tr[\rho_{\lambda}(g^{-1})\widehat{m}_{\alpha}(\lambda)\widehat{m}_{\beta}(\lambda)]$
(3.24) $=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\dim V_{\lambda}\cdot Tr\left[\frac{\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})}{\dim V_{\lambda}^{2}}\rho_{\lambda}(g^{-1})\right]$
(3.25) $=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\frac{1}{\dim V_{\lambda}}\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\chi_{\lambda}(g^{-1})$
@@
The induced density on $\gamma$ is given by the density (3.22) multiplied by the Jacobian of diagonalization map $g\mapsto V\mathrm{e}^{2i\pi\gamma}V^{*}$ with $\gamma=(\gamma_{1},\dots,\gamma_{n})$ such that $\sum\gamma_{i}\in\mathbb{Z}$. Since this Jacobian is $\frac{|\Delta(\mathrm{e}^{2i\pi\gamma})|^{2}}{(2\pi)^{n-1}n!}$, see [7, eq. (17)], we obtain the desired expression. ∎

### v2

@@ 2. Notations and statement of the result
where $\Delta(e^{2i\pi\theta})=2^{n(n-1)/2}\prod_{i<j}\sin\left(\pi(\theta_{i}-\theta_{j})\right)$ for $\theta\in\mathcal{H}$ and $Vol_{g}$ denotes the volume with respect to the Lebesgue measure on $\mathbb{R}^{Supp(g)\setminus\partial R_{d,n}}$.
Note that the case $n=2$ admits explicit formulas which do not need such a machinery. Numerical experiments for $n=3$ suggest that there are $\alpha,\beta\in\mathcal{H}_{reg}$ for which any regular labeling $g$ yields a non-empty polytope $P^{g}_{\alpha,\beta,\gamma}$ for some $\gamma\in\mathcal{H}_{reg}$. However, for a fixed triple $(\alpha,\beta,\gamma)\in\mathcal{H}_{reg}$ there seems to be generically only a strict subset of $\{P^{g}_{\alpha,\beta,\gamma}\}_{g\text{ regular}}$ which are not empty and do contribute. Finally, remark that we only considered the case of regular conjugacy classes to ensure the existence of a density for the convolution product. Such a hypothesis is regularly assumed (see for example [36, 27]). However, we expect similar results to hold in cases where less than $n/2$ coordinates of $\mathcal{H}$ are equal, in which case the convolution product is still expected to have a density.
Volume of moduli spaces of flat $SU_{n}$-connections on the sphere:
The main application of the previous result is the computation of the volume of moduli spaces of flat $SU_{n}$-connections on the three-punctured sphere. Computing such volume is an important task in the study of the Yang-Mills measure on Riemann surfaces in the small surface limit [13], and it has been shown in [37, 27] that this computation for arbitrary compact Riemann surfaces can be reduced to case of the three-punctured sphere by a sewing phenomenon. A similar inductive procedure is used in [28] to reduce the volume problem for the moduli space of curves to the genus zero case.
Denote by $\Sigma_{0}^{3}$ the sphere with three generic marked points $a,b,c$ removed. We then denote by $\mathcal{M}(\Sigma_{0}^{3},\alpha,\beta,\gamma)$ the moduli space of flat $SU_{n}$-valued connections on $\Sigma_{0}^{3}$ for which the holonomies around $a,b,c$ respectively belong to $\mathcal{O}_{\alpha},\mathcal{O}_{\beta}$ and $\mathcal{O}_{\gamma}$. In the specific case of the punctured sphere, this moduli space can be alternatively described as
@@
where $SU_{n}$ acts diagonally by conjugation, see [1], see also [27, Section 3] for a general introduction to $SU_{n}$-valued flat connections and a proof of the latter equality. As a corollary of Theorem 2.5, we thus get an expression of the volume of $\mathcal{M}(\Sigma_{0}^{3},\alpha,\beta,\gamma)$ as a sum of volumes of explicit polytopes.
Corollary 2.6 (Volume of flat $SU_{n}$-connections on the sphere).
Let $n\geq 3$ and consider the canonical volume form on $SU_{n}$. For $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ such that $|\alpha|_{1},\,|\beta|_{1},\,|\gamma|_{1}\in\mathbb{N}$, then $Vol\left[(\mathcal{M}(\Sigma_{0}^{3},\alpha,\beta,\gamma)\right]\not=0$ only if $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}+\sum_{i=1}^{n}\gamma_{i}=n+d$ for some $d\in\mathbb{N}$, in which case
@@
where $\widetilde{\gamma}=(1-\gamma_{n},\ldots,1-\gamma_{1})$ and the polytopes $P^{g}_{\alpha,\beta,\widetilde{\gamma}}$ are defined in Theorem 2.5.
Note that the choice of normalization for the volume of $SU_{n}$ slightly differs from the one used in [36] for numerical applications. As a consequence of this corollary, the volume is a piecewise polynomial in $\alpha,\beta,\gamma$, up to the normalization factor coming from the volume of the conjugacy classes. Such a phenomenon, which is a reflect of the underlying symplectic structure, had already been observed in [27]. A same phenomenon occurs in the co-adjoint case, see [8, 11] and in the study of moduli spaces of curves, [28].
@@
Let us give the structure of the paper while sketching the proof of our main statement. The first step of the prooF is the semi-classical approximation of the density of the convolution product by a limit of quantum Littlewood-Richardson coefficients. Such approximation scheme is done in Section 3. All the work of the remaining part of the manuscript consists in turning known expressions for the quantum Littlewood-Richardson coefficients into integers points counting in convex bodies, for which the convergence towards volumes of polytopeS is straightforward, see [19]. Section 4 introduces the puzzle expression for those coefficients obtained in [6] and gives a first simplification of the puzzle formulation by only keeping the position of certain pieces of the puzzles. It is then deduced in Section 5 an expression of the coefficients as the counting of integers points in a family of convex polytopes indexed by certain two-colored tilings which are reminiscent of Figure 4, see Theorem 4.5 and Corollary 5.4. Up to this point, those polytopes are degenerated and non-rational polytopes in a higher dimensionAL space, which prevents any proper asymptotic counting in the semi-classical limit (a similar problematic situation already occurred with Berenstein-Zelevinsky polytopes in the co-adjoint case, leading to the hive formulation of Knutson and Tao, see [19]). By a combinatorial work on the underlying two-colored tiling, we give in Section 6 a parametrization of the integer points of those polytopes in terms of integer points of genuine convex bodies. Remark that the results of Section 5 hold more generally for any coefficient of the two-step flag variety, a fact which is not true anymore from Section 6. The asymptotic counting of integers points in convex bodies is then much more tractable, and the conclusion of the proof of Theorem 2.5 and Corollary 2.6 is done in Section 7.
3. Density formula via the quantum cohomology of the Grassmannians
Let $n\geq 1$ and consider $\alpha,\beta\in\mathcal{H}_{reg}^{2}$ : $\alpha=(\alpha_{1},\dots,\alpha_{n})$, $\beta=(\beta_{1},\dots,\beta_{n})$ where
@@ 3. Density formula via the quantum cohomology of the Grassmannians
Up to multiplication by the center of $U(n)$, suppose furthermore without loss of generality that
@@
for some $k,k^{\prime}\in\mathbb{Z}$. Let $A=U\mathrm{e}^{2i\pi\alpha}U^{*},B=V\mathrm{e}^{2i\pi\beta}V^{*}$, where $U,V$ are independent Haar distributed matrices on $U(n)$. Remark that $A$ and $B$ are respectively uniformly distributed on the conjugacy classes $\mathcal{O}(\alpha)$ and $\mathcal{O}(\beta)$, which lie in $SU(n)\subset U(n)$. The goal of this section, see Theorem 3.8, is to give a simple proof of the density formula (3.26) linking the probability $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ that $AB\in\mathcal{O}(\gamma)$ for $\gamma\in\mathcal{H}$ to the structure constants of the quantum cohomology of Grassmannians defined in Section 3.2. Such a semi-classical convergence had been already suggested and proven several times in different forms (see [36] for a similar approach with fusion coefficient and [9] for a convergence in distribution). In Section 3.1, we recall in Proposition 3.1 a classical expression of the density in terms of characters of irreducible representations of $SU(n)$. In Section 3.2, we link the density of Proposition 3.1 to the structure constants and derive Theorem 3.8.
@@ 3.1. A first density formula
This part aims at recalling a proof of the formula (3.3) which gives the value of $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ as an infinite sum of characters. A similar treatment of the convolution of orbit measures in the general context of Lie algebras can be found in [9, Sec. 7].
@@
for the push-forward of $\mathrm{d}g$ by $\varphi_{\theta}$. The measure $m_{\theta}$ is a measure on $\mathcal{O}(\theta)$ called the orbital measure. For any function $f:\mathcal{O}(\theta)\rightarrow\mathbb{R}$,
@@
Recall that the irreducible representations of the compact group $SU(n)$ are parameterized by $\lambda\in\mathbb{Z}_{\geq 0}^{n-1}$ and we denote by $(\rho_{\lambda},V_{\lambda})$ the corresponding representation where $\rho_{\lambda}:SU(n)\rightarrow V_{\lambda}$ and $\chi_{\lambda}:End_{V_{\lambda}}\rightarrow\mathbb{C},x\mapsto Tr[x]$ is the associated character.
@@
Let $(\alpha,\beta)\in\mathcal{H}_{reg}^{2}$ and let $A,B\in\mathcal{O}(\alpha)\times\mathcal{O}(\beta)$ be two independent random variables sampled from $m_{\alpha}$ and $m_{\beta}$ respectively. Let $C=AB\in\mathcal{O}(\gamma)$ for some random $\gamma\in[0,1[^{n}/S_{n}$. The density of $\gamma=\gamma_{1}>\dots>\gamma_{n}\geq 0$ is given by the absolute convergent series
(3.3) $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]=\frac{|\Delta(\mathrm{e}^{2i\pi\gamma})|^{2}}{(2\pi)^{n-1}n!}\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\frac{1}{\dim V_{\lambda}}\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\chi_{\lambda}(\mathrm{e}^{-2i\pi\gamma}).$
Another expression of the density (3.3) is given in (3.24). The rest of this section is devoted to the proof of Proposition 3.1.
@@
(3.4) $\widehat{m}:\lambda\in\mathbb{Z}_{\geq 0}^{n-1}\mapsto\widehat{m}(\lambda)=\int_{SU(n)}\rho_{\lambda}(g)\mathrm{d}m(g)\in\mathrm{End}_{V_{\lambda}}.$
In the case where $m=m_{\theta}$, the expression $\widehat{m_{\theta}}(\lambda)$ is also known as the spherical transform introduced in [40, eq. (56)].
@@
For any $\lambda\in\mathbb{Z}_{\geq 0}^{n-1}$ and $g\in SU(n)$, since the Haar measure is invariant by translation,
@@
[…]
Hence, $\widehat{m_{\theta}}(\lambda)$ is a morphism of the irreducible representation $\rho_{\lambda}$ and thus $\widehat{m_{\theta}}(\lambda)=c\cdot\mathrm{id}_{V_{\lambda}}$ for some $c\in\mathbb{C}$. One computes the value of $c$ by taking the trace which gives
@@
Let $m,m^{\prime}$ be two measures on $SU(n)$. Let $m\boxtimes m^{\prime}$ be the product measure on $SU(n)\times SU(n)$. Define $mult:SU(n)\times SU(n)\rightarrow SU(n)$ to be the multiplication on $SU(n)$ : $mult(g_{1},g_{2})=g_{1}g_{2}$. The convolution of $m$ and $m^{\prime}$, denoted by $m*m^{\prime}$, is defined as
(3.7) $m*m^{\prime}=mult\#(m\boxtimes m^{\prime}),$
@@
By Definition 3.4, the measure $m_{\alpha,\beta}$ is the law on $SU(n)$ of $C=A\cdot B$ where $A$ and $B$ are sampled from measures $\mu_{\alpha}$ and $\mu_{\beta}$ on $\mathcal{O}(\alpha)$ and $\mathcal{O}(\beta)$ respectively. Recall that for two measures $m,m^{\prime}$ on $SU(n)$,
@@
Let $f:\lambda\in\mathbb{Z}_{\geq 0}^{n-1}\mapsto f(\lambda)\in End_{V_{\lambda}}$ be a function such that
(3.10) $\left\lVert f\right\rVert^{2}=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\dim V_{\lambda}\cdot Tr[f(\lambda)f^{*}(\lambda)]<\infty.$
@@
(3.12) $g$ $\mapsto\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\dim V_{\lambda}\cdot Tr[\rho_{\lambda}(g^{-1})f(\lambda)].$
@@
For $(\alpha,\beta)\in\mathcal{H}_{reg}^{2}$,
(3.13) $\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\dim(V_{\lambda})Tr[\widehat{m}_{\alpha,\beta}(\lambda)\widehat{m}_{\alpha,\beta}(\lambda)^{*}]<\infty.$
@@
(3.14) $\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\dim(V_{\lambda})Tr[\widehat{m}_{\alpha,\beta}(\lambda)\widehat{m}_{\alpha,\beta}(\lambda)^{*}]=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\frac{1}{\dim(V_{\lambda})^{2}}\left|\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\right|^{2}\left|\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\right|^{2}.$
Using Weyl’s character formula [8, eq. (21)],
@@
where $\lambda^{\prime}=(\lambda_{1},\ldots,\lambda_{n-1},0)+\rho$ with $\rho=(n-1,\ldots,0)$ and where $\Delta(x_{1},\dots,x_{n})=\prod_{1\leq i<j\leq n}(x_{i}-x_{j})$ is the Vandermonde determinant. Recall that by assumption, $\alpha_{i}\neq\alpha_{j}$ for $i\neq j$ and the same holds for $\beta$, so that the expressions $\Delta(\mathrm{e}^{2i\pi\alpha})$ and $\Delta(\mathrm{e}^{2i\pi\beta})$ are well-defined. The previous sum becomes
@@
which can be found in [12, Cor. 11.2.5] and where $sf(n)=\prod_{1\leq j\leq n}j!$, it suffices to show that
(3.19) $V_{n}=\sum_{\lambda_{1}>\dots>\lambda_{n}=0}\frac{1}{\Delta(\lambda)^{2}}$
converges for $n\geq 2$. One has that $V_{2}=\sum_{k\geq 1}k^{-2}<\infty$. Let us write
$\sum_{\lambda_{1}>\dots>\lambda_{n}=0}\frac{1}{\Delta(\lambda)^{2}}$ $=\sum_{\lambda_{1}>\dots>\lambda_{n}=0}\prod_{1\leq i<j\leq n}(\lambda_{i}-\lambda_{j})^{-2}$
$=\sum_{\lambda_{2}>\dots>\lambda_{n}=0}\left[\sum_{\lambda_{1}>\lambda_{2}}\prod_{2\leq j\leq n}(\lambda_{1}-\lambda_{j})^{-2}\right]\prod_{2\leq i<j\leq n}(\lambda_{i}-\lambda_{j})^{-2}$
$\leq\sum_{\lambda_{2}>\dots>\lambda_{n}=0}\left[\sum_{\lambda_{1}>\lambda_{2}}(\lambda_{1}-\lambda_{2})^{-2(n-1)}\right]\prod_{2\leq i<j\leq n}(\lambda_{i}-\lambda_{j})^{-2}$
the innermost sum is bounded by $\sum_{k\geq 1}k^{-2(n-1)}\leq\sum_{k\geq 1}k^{-2}=V_{2}$ for $n\geq 2$. Thus,
$V_{n}\leq V_{2}V_{n-1}$
so that for $n\geq 2$, $V_{n}\leq(V_{2})^{n-1}$ which proves the convergence.
@@
Let $(\alpha,\beta)\in\mathcal{H}_{reg}^{2}$ and $g\in SU(n)$. Then,
(3.20) $(\widehat{m}_{\alpha,\beta})^{\lor}(g)=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\frac{1}{\dim V_{\lambda}}\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\chi_{\lambda}(g^{-1}),$
where the sum converges in $L^{2}(SU_{n})$.
@@
(3.21) $(\widehat{m}_{\alpha,\beta})^{\lor}(g)$ $=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\dim V_{\lambda}\cdot Tr[\rho_{\lambda}(g^{-1})\widehat{m}_{\alpha}(\lambda)\widehat{m}_{\beta}(\lambda)]$
(3.22) $=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\dim V_{\lambda}\cdot Tr\left[\frac{\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})}{\dim V_{\lambda}^{2}}\rho_{\lambda}(g^{-1})\right]$
(3.23) $=\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\frac{1}{\dim V_{\lambda}}\chi_{\lambda}(\mathrm{e}^{2i\pi\alpha})\chi_{\lambda}(\mathrm{e}^{2i\pi\beta})\chi_{\lambda}(g^{-1}).$
@@
The induced density on $\gamma$ is given by the density (3.20) multiplied by the Jacobian of the diagonalization map $g\mapsto V\mathrm{e}^{2i\pi\gamma}V^{*}$ with $\gamma=(\gamma_{1},\dots,\gamma_{n})$ such that $\sum\gamma_{i}\in\mathbb{Z}$. Since this Jacobian is $\frac{|\Delta(\mathrm{e}^{2i\pi\gamma})|^{2}}{(2\pi)^{n-1}n!}$, see [12, Thm 11.2.1], we obtain the desired expression. ∎